OpenStudy (anonymous):

Please help me. 6+x over 3=1

OpenStudy (anonymous):

0k

OpenStudy (anonymous):

\[\frac{6+x}{3}=1\] multiply by LCD (3) \[3*\frac{6+x}{3}=3*1\] \[6+x=3\] subtract 6 from both sides and you should get the answer

OpenStudy (anonymous):

-3...

OpenStudy (anonymous):



OpenStudy (anonymous):